# Delving Deeper Into Astromorphic Transformers

### 18 Dec 2023 v1

@@ Abstract
Introduction
@@
Astrocytes, a type of glial cell, play a critical role in brain function encompassing various processes such as homeostasis, metabolism, and synaptic regulation (Perez-Catalan, Doe, and Ackerman 2021). Astrocytes detect and regulate synaptic activity in the tripartite synapse through interactions with pre and postsynaptic neurons. Investigating their impact on neural computation is currently an active research field in neuroscience and underscores the critical need to move beyond the neuro-synaptic perspective of current Artificial Intelligence (AI) systems. Recent experimental findings on neuron-astrocyte interactions and modulation have led to significant progress in computational neuroscience, enabling the development of models that incorporate neuron-astrocyte interactions within neural networks (Manninen, Aćimović, and Linne 2023; Wakida et al. 2020). Astrocytes have been found to modulate bursting in neural circuitry through the release of gliotransmitters, which have an impact on neuronal excitability and synaptic plasticity (Durkee et al. 2021; Lenk et al. 2020). Astrocytes possess the ability to encode information through calcium signaling and regulate information processing, thereby actively engaging in neural computation at the tripartite synapse level. Additionally, astrocytes possess inherent capacity as memory components (Tsybina et al. 2022; Gordleeva et al. 2022) and plasticity regulators that are capable of facilitating local sequential learning (Gibbs, Hutchinson, and Hertz 2008; Alberini et al. 2018).
@@
Preliminary attempts at incorporating the critical role of astrocytes in brain-inspired neuromorphic computing has started in earnest. Neuromorphic systems (Sengupta et al. 2019) are bio-plausible AI models that attempt to emulate neuronal and synaptic operations with a higher degree of bio-fidelity to tap into the huge energy efficiency of the brain. Such systems are temporal and event-driven where computation occurs through transmission of sparse spiking events between neurons - thereby offering the potential of orders of magnitude power efficiency when implemented on neuromorphic platforms (Davies et al. 2021). However, astromorphic computing encapsulating features of astrocytic regulation in neuromorphic systems still remain in infancy with works primarily targeting integration of astrocyte functionality with Spike-Timing-Dependent-Plasticity (STDP) mechanisms (Liu et al. 2018; Rastogi et al. 2021a; Han, Islam, and Sengupta 2022; Liu et al. 2017) - an unsupervised learning framework that remains limited to shallow networks and simple tasks (Diehl and Cook 2015). Other attempts also suffer from the pressing issue of scalability to complex tasks (Han, Pan, and Liu 2023).
@@
This work is based on a recent study (Kozachkov, Kastanenka, and Krotov 2023) that has shown promise at emulating the self-attention mechanism in transformers (Vaswani et al. 2017) using a simple two layered network involving neuron-astrocyte interactions. The ability to design Astromorphic Transformers is therefore highly relevant in the current era of Large Language Models (LLMs), such as BERT (Devlin et al. 2018) and GPT (Radford et al. 2018) - paving the pathway for designing explainable transformers from a neuroscience perspective as well as addressing scaling challenges for utilizing the computational role of astrocytes in neuromorphic computing. However, as we illustrate in this work, prior work has neglected various key aspects of bi-directional astrocytic signalling, feedback and non-linear temporal behavior - all of which are key to address the fundamental mapping of self-attention mechanism to astrocyte-neuron interactions along with providing insights on whether such intrinsic non-linear dynamics and plasticity modulation enable better learning capabilities from the application side. The distinguishing feature of this work therefore lies in the neuroscience-algorithm-application co-design perspective seeking to answer the following questions: (i) Theoretical neuroscience: What aspects of astrocyte functionality should be included in neural architectures to better explain self-attention mechanism in transformers? (ii) Algorithm: Can we map key algorithmic modules like relative positional information encoding of tokens to astrocytic interactions by drawing qualitative inspiration from theoretical neuroscience? (iii) Application: What is the impact of incorporating such dynamic temporal bio-plausible transformers from the machine learning application perspective?
Related Works and Main Contributions
@@
Understanding the role of astrocytes in neuromorphic computing, especially from the algorithm perspective, is relatively nascent. While hardware implementation of astrocyte functionality in CMOS (Ranjbar and Amiri 2015; Karimi et al. 2018) as well as post-CMOS technologies (Garg, Yang, and Sengupta 2021) have been explored, algorithm level works exploiting the role of astrocytes and leveraging their computational role has largely remain limited to self-repair functionality of faulty neural network hardware (Wade et al. 2012a; Liu et al. 2018; Rastogi et al. 2021b; Han, Islam, and Sengupta 2022). This is mainly inspired by the fact that astrocytes, through their synaptic plasticity regulation and feedback mechanisms, enable neurons to fire at a baseline ideal frequency. This property can be utilized to mitigate neuromorphic hardware non-idealities with stuck-at-faults and drift effects (Han, Islam, and Sengupta 2022). However, such approaches are still limited to single layered unsupervised networks where astrocyte functionality is integrated with STDP learning mechanisms.
Recent studies have also examined the influence of astrocytes on learning (Alberini et al. 2018), working memory (Tsybina et al. 2022), energy minimization in neural networks (Fountain and Merkel 2022), robotic locomotion control (Polykretis, Tang, and Michmizos 2020), structure learning (Han, Pan, and Liu 2023), among others.
@@
Building upon recent works (Kozachkov, Kastanenka, and Krotov 2023) at emulating self-attention module in transformers through astrocyte-neuron interactions, we develop a computational framework that encapsulates key aspects of temporal non-linearities and feedback enabled by astrocytes and underscore the importance of incorporating such bio-plausible effects through a comprehensive evaluation of Astromorphic Transformers on a benchmark evaluation set. Our work expands the computational capabilities of astrocyte-neuron network architectures equipped with plasticity, for the first time, to a broad category of sequential tasks.
Results
Transformers
@@ Results
Vanilla transformers (Vaswani et al. 2017) for classification tasks usually consist of an encoder block processing the provided key$(K)$, query$(Q)$ and value$(V)$ generated from the input token embeddings $X$ (words in language tasks, pixels in vision tasks, etc.). Appendix A in the supplementary material illustrates the conventional configuration of a transformer, featuring a block referred to as “Multi-Head Attention” that produces the contextual association among the provided inputs. The Multi-Head Attention block in transformers uses self-attention mechanisms for multiple input heads, allowing the model to allocate varying importance to individual tokens, focusing on salient features for precise output in specific contexts.
@@
The transformer implements auto-regressive attention by computing the dot product of key and query, and then multiplying the resulting scalar with the corresponding value derived from the input token. Implementing the softmax function (a type of distribution function) atop the dot product of key and query (to enforce attention over the tokens) increases the order of complexity, a predicament that can be alleviated by the use of linearized attention (Schlag, Irie, and Schmidhuber 2021). Linearized attention exploits the distributive property of matrix multiplication resulting in lower complexity for self-attention computation. More details can be found in Appendix B of the supplementary material. From the linear attention-based transformer (linear transformer) perspective (Katharopoulos et al. 2020), the transformer equations can be defined in the following form:
@@
From Equations 2 and 3, we can see that the output of self-attention is $Y$, which undergoes layer normalization after being augmented with residual connections from the input layer. The transformer determines the final output probabilities, denoted by $Z$. There is a $\phi(.)$ feature map that corresponds to a positive similarity function. When two such functions are multiplied, the normalized dot product maps to a distribution function (Chaudhari et al. 2021) (i.e. softmax). According to (Katharopoulos et al. 2020), $\phi(.)$ can be defined by Equation 4. Here, the activation function $elu(x)$, which stands for exponential linear unit, is utilized for efficient and precise learning (Clevert, Unterthiner, and Hochreiter 2015).
@@
Tripartite Synapse Dynamics
@@
Astrocyte dynamics:
The tripartite synapse’s dynamics are based on Li and Rinzel $Ca^{2+}$ dynamics (Li and Rinzel 1994), gatekeeper model (Volman, Ben-Jacob, and Levine 2007), Nadkarni and Jung model (Nadkarni and Jung 2004, 2007), and bidirectional study on astrocytic glutamate binding to postsynaptic neurons (AN model) (Wade et al. 2011).
@@
According to the biophysical models, it is observed that an action potential (AP) is generated by the presynaptic axon through the release of neurotransmitters. Neurotransmitters bind to postsynaptic dendrites, causing depolarization. This leads to the release of 2-Arachidonylglycerol (2-AG) from postsynaptic neurons to type 1 Cannabinoid Receptors (CB1Rs) on the presynaptic terminal. This process is called Depolarization-induced Suppression of Excitation ($DSE$) and depresses the continuous release of neurotransmitters by inhibiting presynaptic neurons (Alger 2002). Neurotransmitter release also involves binding 2-AG to CB1Rs on an astrocyte, enveloping the synapse and increasing Inositol 1,4,5-trisphosphate ($IP_{3}$) levels. $IP_{3}$ facilitates the release of calcium ions into the cytoplasm, thereby, increasing calcium concentration. Three currents, $J_{channel}$, $J_{leak}$, and $J_{pump}$, are introduced to model the rate of change of calcium concentration inside the astrocyte (Li and Rinzel 1994). When calcium concentration crosses a threshold, astrocytes discharge glutamate, which binds to presynaptic group I metabotropic Glutamate Receptors (mGluRs), creating a process called eSP (Indirect Feedback) from the astrocyte to the presynaptic neuron (Navarrete and Araque 2010). The comprehensive description of the computational model (Equation 5 introduces the basic equation of the model) can be found in the works of (Wade et al. 2012b, 2011; De Pittà et al. 2009).
@@
$I_{syn}^{i}$ is the intrasynaptic current which varies linearly with the amount of neurotransmitters present in that synapse at any time, $t$ ($I_{syn}^{i}=A_{se}y_{i}(t)$ where, $A_{se}$ and $y_{i}(t)$ represent the absolute synaptic efficacy and the amount of neurotransmitter discharged by the synapse at a given time $t$, respectively). Thus, $I_{syn}$ depends on the direct feedback (corresponds to $DSE$) as direct feedback is modulated by the amount of neurotransmitters between the pre and postsynaptic neuron. On the other hand, the extrasynaptic Slow Inward Current $(SIC(t))$ is generated when astrocyte-released glutamate interacts with extrasynaptic NMDARs (located in postsynaptic neuron) (Haydon and Carmignoto 2006). To elaborate, the activation of astrocytes leads to the release of glutamate from these cells. Upon its release, glutamate interacts with NMDARs located on the postsynaptic neuron, resulting in the emergence of a distinct form of current known as Slow Inward Current $(SIC(t))$ (Wade et al. 2011). The details of the mathematical modeling of $SIC(t)$ can be found in (Wade et al. 2011).
@@
Neural Network Emulating Self-Attention
Self-attention is a fundamental component of all transformer models. Based on prior works of mimicking self-attention by astrocytes (Kozachkov, Kastanenka, and Krotov 2023), we adopt a network topology as shown in Fig. 2. The neural network architecture comprises three distinct layers, namely the Input layer, Hidden layer, and Output layer. There exists two distinct operational modes for the aforementioned neural network: (i) write mode and (ii) read mode. Upon presentation of the tokens ($x_{t}\in R^{1\times d}$: which may include word embeddings and image embeddings) to the input layer, the write mode is initially employed, followed by the read mode to generate the outputs. In write mode, we encode information into the network by means of keys and values, whereas, in read mode, we retrieve the information by providing queries.
@@
Mathematical Formulation of Plasticity
(i) Hebbian Plasticity:
Hebbian plasticity is characterized by the sequential adjustment of synaptic weights in accordance with the interconnection pattern between two neurons (pre and postsynaptic neurons) at the synapse. Previous studies (Kozachkov, Kastanenka, and Krotov 2023) have exclusively focused on the plasticity occurring between two neurons neglecting non-linearities and the plasticity that may occur between astrocytes and postsynaptic neurons, which is also involved in the learning process.
@@
Astrocytes have a significant impact on the Hebbian weight due to their inherent non-linearity. A recent study by (Han, Pan, and Liu 2023) introduced an AstroNet model that optimizes neural network connections through temporal and global regulation mechanisms. The authors found that astrocytes introduce non-linear effects, which compound the pre-existing nonlinearity of neurons. The concept of Hebbian Plasticity is associated with temporal and global regulation, where the current weight is influenced by past neuron weights and astrocytic weights. To map the effect of non-linearity, we consider that the summation of neuron and astrocyte Hebbian weights undergoes non-linear activation, ultimately generating the Hebbian weight ($H$). As such, we propose the integration of a sigmoid ($\sigma$) nonlinearity into the learning process. Thus, the weight $H$ exhibits the collective Hebbian plasticity of a tripartite synapse, as elucidated by Equation 11, where $K$ and $V$ correspond to keys and values of all the $N$ tokens in a single sequence (in matrix form: $K\in R^{N\times m}$, $V\in R^{N\times d}$). Assuming initial $H_{neuron,0}$ and $H_{astro,0}$ from Equations 9 and 10 are both zero, we can define the Hebbian weight in its matrix form ($H$) as,
@@
Next, we consider formulation of the astrocytic activity parameter, $W_{astro}$. We note that astrocytic $Ca^{2+}$ dynamics is a relatively slower temporal process than the streaming rate of the input tokens. Therefore, it is expected that the astrocytic activity should encode relative information among successive tokens. In this context, we derive inspiration from algorithmic formulations of relative positional information of tokens, popularly used in Vision transformers (Shaw, Uszkoreit, and Vaswani 2018; Ramachandran et al. 2019), and map it to the activity parameter $W_{astro}$. The astrocytic response parameter is therefore defined to use the edge of the tokens to capture relative position differences between input elements (Shaw, Uszkoreit, and Vaswani 2018). The edge of the tokens considers the pairwise relationships between input elements in the sense that the input is modeled as a labeled, directed, fully-connected graph. The edges pertaining to relative positional information are learnt using the temporal regulation property of astrocytes through the activity parameter, $W_{astro}$. Consequently, $H_{astro}$ can be considered as a representation of acquired positional knowledge derived from the input tokens. $W_{astro}$ can be modeled using Equation 12 where, $a_{ij}$ is the edge between input tokens $x_{i}$ and $x_{j}$.

### 09 Jan 2024 v2

@@ Abstract
Index Terms:
Astromorphic Transformer, Tripartite Synapse, Hebbian Plasticity, Presynaptic Plasticity, Self-Attention
I Introduction
@@
Astrocytes, a type of glial cell, play a critical role in brain function encompassing various processes such as homeostasis, metabolism, and synaptic regulation [1]. Astrocytes detect and regulate synaptic activity in the tripartite synapse through interactions with pre and postsynaptic neurons. Investigating their impact on neural computation is currently an active research field in neuroscience and underscores the critical need to move beyond the neuro-synaptic perspective of current Artificial Intelligence (AI) systems. Recent experimental findings on neuron-astrocyte interactions and modulation have led to significant progress in computational neuroscience, enabling the development of models that incorporate neuron-astrocyte interactions within neural networks [2, 3]. Astrocytes have been found to modulate bursting in neural circuitry through the release of gliotransmitters, which have an impact on neuronal excitability and synaptic plasticity [4, 5]. Astrocytes possess the ability to encode information through calcium signaling and regulate information processing, thereby actively engaging in neural computation at the tripartite synapse level. Additionally, astrocytes possess inherent capacity as memory components [6, 7] and plasticity regulators that are capable of facilitating local sequential learning [8, 9].
@@
Preliminary attempts at incorporating the critical role of astrocytes in brain-inspired neuromorphic computing has started in earnest. Neuromorphic systems [10] are bio-plausible AI models that attempt to emulate neuronal and synaptic operations with a higher degree of bio-fidelity to tap into the huge energy efficiency of the brain. Such systems are temporal and event-driven where computation occurs through transmission of sparse spiking events between neurons - thereby offering the potential of orders of magnitude power efficiency when implemented on neuromorphic platforms [11]. However, astromorphic computing encapsulating features of astrocytic regulation in neuromorphic systems still remain in infancy with works primarily targeting integration of astrocyte functionality with Spike-Timing-Dependent-Plasticity (STDP) mechanisms [12, 13, 14, 15] - an unsupervised learning framework that remains limited to shallow networks and simple tasks [16]. Other attempts also suffer from the pressing issue of scalability to complex tasks [17].
@@
This work is based on a recent study [18] that has shown promise at emulating the self-attention mechanism in transformers [19] using a simple two layered network involving neuron-astrocyte interactions. The ability to design Astromorphic Transformers is therefore highly relevant in the current era of Large Language Models (LLMs), such as BERT [20] and GPT [21] - paving the pathway for designing explainable transformers from a neuroscience perspective as well as addressing scaling challenges for utilizing the computational role of astrocytes in neuromorphic computing. However, as we illustrate in this work, prior work has neglected various key aspects of bi-directional astrocytic signalling, feedback and non-linear temporal behavior - all of which are key to address the fundamental mapping of self-attention mechanism to astrocyte-neuron interactions along with providing insights on whether such intrinsic non-linear dynamics and plasticity modulation enable better learning capabilities from the application side. The distinguishing feature of this work therefore lies in the neuroscience-algorithm-application co-design perspective seeking to answer the following questions: (i) Theoretical neuroscience: What aspects of astrocyte functionality should be included in neural architectures to better explain self-attention mechanism in transformers? (ii) Algorithm: Can we map key algorithmic modules like relative positional information encoding of tokens to astrocytic interactions by drawing qualitative inspiration from theoretical neuroscience? (iii) Application: What is the impact of incorporating such dynamic temporal bio-plausible transformers from the machine learning application perspective?
II Related Works and Main Contributions
@@
Understanding the role of astrocytes in neuromorphic computing, especially from the algorithm perspective, is relatively nascent. While hardware implementation of astrocyte functionality in CMOS [22, 23] as well as post-CMOS technologies [24] have been explored, algorithm level works exploiting the role of astrocytes and leveraging their computational role has largely remain limited to self-repair functionality of faulty neural network hardware [25, 12, 26, 14]. This is mainly inspired by the fact that astrocytes, through their synaptic plasticity regulation and feedback mechanisms, enable neurons to fire at a baseline ideal frequency. This property can be utilized to mitigate neuromorphic hardware non-idealities with stuck-at-faults and drift effects [14]. However, such approaches are still limited to single layered unsupervised networks where astrocyte functionality is integrated with STDP learning mechanisms.
@@
Recent studies have also examined the influence of astrocytes on learning [9], working memory [6], energy minimization in neural networks [27], robotic locomotion control [28], structure learning [17], among others. Building upon recent works [18] at emulating self-attention module in transformers through astrocyte-neuron interactions, we develop a computational framework that encapsulates key aspects of temporal non-linearities and feedback enabled by astrocytes and underscore the importance of incorporating such bio-plausible effects through a comprehensive evaluation of Astromorphic Transformers on a benchmark evaluation set. Our work expands the computational capabilities of astrocyte-neuron network architectures equipped with plasticity, for the first time, to a broad category of sequential tasks.
III Methodology
III-A Transformers
@@
Vanilla transformers [19] for classification tasks usually consist of an encoder block processing the provided key$(K)$, query$(Q)$ and value$(V)$ generated from the input token embeddings $X$ (words in language tasks, pixels in vision tasks, etc.). Appendix A in the supplementary material illustrates the conventional configuration of a transformer, featuring a block referred to as “Multi-Head Attention” that produces the contextual association among the provided inputs. The Multi-Head Attention block in transformers uses self-attention mechanisms for multiple input heads, allowing the model to allocate varying importance to individual tokens, focusing on salient features for precise output in specific contexts.
@@
The transformer implements auto-regressive attention by computing the dot product of key and query, and then multiplying the resulting scalar with the corresponding value derived from the input token. Implementing the softmax function (a type of distribution function) atop the dot product of key and query (to enforce attention over the tokens) increases the order of complexity, a predicament that can be alleviated by the use of linearized attention [29]. Linearized attention exploits the distributive property of matrix multiplication resulting in lower complexity for self-attention computation. More details can be found in Appendix B of the supplementary material. From the linear attention-based transformer (linear transformer) perspective [30], the transformer equations can be defined in the following form:
@@
From Equations 2 and 3, we can see that the output of self-attention is $Y$, which undergoes layer normalization after being augmented with residual connections from the input layer. The transformer determines the final output probabilities, denoted by $Z$. There is a $\phi(.)$ feature map that corresponds to a positive similarity function. When two such functions are multiplied, the normalized dot product maps to a distribution function [31] (i.e. softmax). According to [30], $\phi(.)$ can be defined by Equation 4. Here, the activation function $elu(x)$, which stands for exponential linear unit, is utilized for efficient and precise learning [32].
@@
III-B Tripartite Synapse Dynamics
@@
III-B1 Astrocyte dynamics
The tripartite synapse’s dynamics are based on Li and Rinzel $Ca^{2+}$ dynamics [33], gatekeeper model [34], Nadkarni and Jung model [35, 36], and bidirectional study on astrocytic glutamate binding to postsynaptic neurons (AN model) [37].
@@ III-B1 Astrocyte dynamics
According to the biophysical models, it is observed that an action potential (AP) is generated by the presynaptic axon through the release of neurotransmitters. Neurotransmitters bind to postsynaptic dendrites, causing depolarization. This leads to the release of 2-Arachidonylglycerol (2-AG) from postsynaptic neurons to type 1 Cannabinoid Receptors (CB1Rs) on the presynaptic terminal. This process is called Depolarization-induced Suppression of Excitation ($DSE$) and depresses the continuous release of neurotransmitters by inhibiting presynaptic neurons [38]. Neurotransmitter release also involves binding 2-AG to CB1Rs on an astrocyte, enveloping the synapse and increasing Inositol 1,4,5-trisphosphate ($IP_{3}$) levels. $IP_{3}$ facilitates the release of calcium ions into the cytoplasm, thereby, increasing calcium concentration. Three currents, $J_{channel}$, $J_{leak}$, and $J_{pump}$, are introduced to model the rate of change of calcium concentration inside the astrocyte [33]. When calcium concentration crosses a threshold, astrocytes discharge glutamate, which binds to presynaptic group I metabotropic Glutamate Receptors (mGluRs), creating a process called eSP (Indirect Feedback) from the astrocyte to the presynaptic neuron [39]. The comprehensive description of the computational model (Equation 5 introduces the basic equation of the model) can be found in the works of [40, 37, 41].
@@ III-B2 Neuron dynamics
$I_{syn}^{i}$ is the intrasynaptic current which varies linearly with the amount of neurotransmitters present in that synapse at any time, $t$ ($I_{syn}^{i}=A_{se}y_{i}(t)$ where, $A_{se}$ and $y_{i}(t)$ represent the absolute synaptic efficacy and the amount of neurotransmitter discharged by the synapse at a given time $t$, respectively). Thus, $I_{syn}$ depends on the direct feedback (corresponds to $DSE$) as direct feedback is modulated by the amount of neurotransmitters between the pre and postsynaptic neuron. On the other hand, the extrasynaptic Slow Inward Current $(SIC(t))$ is generated when astrocyte-released glutamate interacts with extrasynaptic NMDARs (located in postsynaptic neuron) [42]. To elaborate, the activation of astrocytes leads to the release of glutamate from these cells. Upon its release, glutamate interacts with NMDARs located on the postsynaptic neuron, resulting in the emergence of a distinct form of current known as Slow Inward Current $(SIC(t))$ [37]. The details of the mathematical modeling of $SIC(t)$ can be found in [37].
@@
III-C Neural Network Emulating Self-Attention
Self-attention is a fundamental component of all transformer models. Based on prior works of mimicking self-attention by astrocytes [18], we adopt a network topology as shown in Fig. 2. The neural network architecture comprises three distinct layers, namely the Input layer, Hidden layer, and Output layer. There exists two distinct operational modes for the aforementioned neural network: (i) write mode and (ii) read mode. Upon presentation of the tokens ($x_{t}\in R^{1\times d}$: which may include word embeddings and image embeddings) to the input layer, the write mode is initially employed, followed by the read mode to generate the outputs. In write mode, we encode information into the network by means of keys and values, whereas, in read mode, we retrieve the information by providing queries.
@@ III-C Neural Network Emulating Self-Attention
III-D Mathematical Formulation of Plasticity
III-D1 Hebbian Plasticity
Hebbian plasticity is characterized by the sequential adjustment of synaptic weights in accordance with the interconnection pattern between two neurons (pre and postsynaptic neurons) at the synapse. Previous studies [18] have exclusively focused on the plasticity occurring between two neurons neglecting non-linearities and the plasticity that may occur between astrocytes and postsynaptic neurons, which is also involved in the learning process.
@@ III-D1 Hebbian Plasticity
Astrocytes have a significant impact on the Hebbian weight due to their inherent non-linearity. A recent study by [17] introduced an AstroNet model that optimizes neural network connections through temporal and global regulation mechanisms. The authors found that astrocytes introduce non-linear effects, which compound the pre-existing nonlinearity of neurons. The concept of Hebbian Plasticity is associated with temporal and global regulation, where the current weight is influenced by past neuron weights and astrocytic weights. To map the effect of non-linearity, we consider that the summation of neuron and astrocyte Hebbian weights undergoes non-linear activation, ultimately generating the Hebbian weight ($H$). As such, we propose the integration of a sigmoid ($\sigma$) nonlinearity into the learning process. Thus, the weight $H$ exhibits the collective Hebbian plasticity of a tripartite synapse, as elucidated by Equation 11, where $K$ and $V$ correspond to keys and values of all the $N$ tokens in a single sequence (in matrix form: $K\in R^{N\times m}$, $V\in R^{N\times d}$). Assuming initial $H_{neuron,0}$ and $H_{astro,0}$ from Equations 9 and 10 are both zero, we can define the Hebbian weight in its matrix form ($H$) as,
@@
Next, we consider formulation of the astrocytic activity parameter, $W_{astro}$. We note that astrocytic $Ca^{2+}$ dynamics is a relatively slower temporal process than the streaming rate of the input tokens. Therefore, it is expected that the astrocytic activity should encode relative information among successive tokens. In this context, we derive inspiration from algorithmic formulations of relative positional information of tokens, popularly used in Vision transformers [43, 44], and map it to the activity parameter $W_{astro}$. The astrocytic response parameter is therefore defined to use the edge of the tokens to capture relative position differences between input elements [43]. The edge of the tokens considers the pairwise relationships between input elements in the sense that the input is modeled as a labeled, directed, fully-connected graph. The edges pertaining to relative positional information are learnt using the temporal regulation property of astrocytes through the activity parameter, $W_{astro}$. Consequently, $H_{astro}$ can be considered as a representation of acquired positional knowledge derived from the input tokens. $W_{astro}$ can be modeled using Equation 12 where, $a_{ij}$ is the edge between input tokens $x_{i}$ and $x_{j}$.